A first-order reaction has a rate constant of 7.5 x10–3 /s. The half-time (half-life) of the reaction is:

A) 0.693 s      B) 92 s      C) 101 s       D) 133 s       E) None of these
Answer #1

We know that

for first order reaction

Half life t1/2 = 0.693 / rate constant k

and we have

k = 7.5 10-3 s-1  

so by putting this value in above formula we get

Half Life

t1/2 = 0.693 / 7.5 10-3 s-1 = 92.4 s

So Our Answer is 92.4 s

From given options there is no value matching to our answer so option E) is correct